To prevent attacks of angina, isosorbide mononitrate is given twice daily typically at doses of 20 mg. It is administered at specific times, such as 8 AM and 3 PM, to prevent nitrate tolerance, which causes the drug to be less effective. How many 10-mg tablets are required for 30 days of treatment?

To prevent attacks of angina, isosorbide dinitrate is given typically at doses of 120 mg/day, divided into three separate doses.
For one week’s regimen, how many 10-mg sublingual tablets of isosorbide dinitrate will be needed?
Answer #1
  • Drug twice daily is 20mg

For 30days = 20×30=600 mg for one month.

So 60 tablets(10mg) is needed

  • 120 mg/ day for one week = 120×7= 840 mg
  • So 84 tablets (10mg) needed